# Automatic Hardware Pragma Insertion in High-Level Synthesis: A Non-Linear Programming Approach

### v1

@@ 2.2. Overview of NLP-DSE
Our primary objective is to efficiently achieve high-quality designs within a limited compilation time budget. To accomplish this we develop a DSE based on a cost model which can find the theoretical optimal configuration in a few seconds/minutes, thanks to its implementation as a Non-Linear Programming (NLP) problem which is then solved using BARON (Tawarmalani and Sahinidis, 2005). Note, however, the in-practice configuration that achieves the best performance may be different: our NLP is unable to model the complexity of the downstream toolchain and issues of resource usage/sharing, actual implementation of the pragmas by Merlin and Vitis, and simple performance bugs in these optimizing compilers may imply a theoretically best design is not implemented as such, making a design space exploration approach still required to achieve a good performance. We have therefore set up a DSE which allows the framework to explore different spaces of parallelism as well as different configuration spaces by constraining the level of parallelism. Our model is presented in Sec. 3, and we prove it is a performance lower bound in Sec. 4, an important feature to be able to prune designs during the search without the risk of losing performance. The associated NLP formulation is given in Sec. 5.
@@
To make our approach feasible, and in particular maintain a sufficient level of accuracy in the analytical models, we target regular, loop-based computations, that is affine programs (Feautrier, 1988). Such programs have a static control-flow, that is every operation executed, and a branch taken at run-time, can be exactly modeled at the compile-time using polyhedral structures (Feautrier, 1992). This restriction improves accuracy of loop-based analytical model by providing them with exact information about the loop trip counts and dependencies.
It is important to note that our DSE process, presented in Sec. 6, fundamentally differs from AutoDSE. We first evaluate configurations that have the lowest theoretical latency, and hence typically a high level of parallelism. However AutoDSE starts from a configuration with no pragmas, which is typically feasible in terms of resource usage, and progressively increases the pragmas until resources are exceeded. Instead our NLP-DSE begins with a design featuring the maximum achievable parallelism, and then gradually reduces the level of parallelism. We illustrate this with Table 1 with the final performance and time-to-solution achieved by NLP-DSE.
As presented in Sec. 7, compared to AutoDSE, our framework locates designs with similar or higher throughput for 46 out of 47 benchmarks evaluated, with a performance improvement of 5.69x and 17.24x on average in terms of the DSE time and design throughput respectively. Only 1 design having a 7% lower throughput. In all 47 benchmarks, the time-to-solution of NLP-DSE is significantly less than AutoDSE, up to 30x faster.
@@ 3.1. Program Representation
We employ the code below as a running example with the pragma above the loops as AMD/Xilinx Merlin and OpenMP. For presentation simplicity, we assume each loop iterator in the program region has been renamed to a unique name, so that we can uniquely identify loops in the code by their iterator name.
@@
6 for(j2=lbj2(i);j2<=ubj2(i);j2++){S2(i,j2);S3(i,j2);}}
The summary AST for this loop is simply built by creating one node per for loop and one per statement Sx, the body of a loop is made of loops and/or statements, and their nodes are children of said loop in the tree, listed in their syntactic order. For example, above, utilizing a constructor notation, it gives: $Loop_{i}(Loop_{j1}(S1),Loop_{j2}(S2,S3))$. Then, a simple rewrites of this tree using loop properties and composition operators will lead to the proposed analytical model, as outlined below.
We now describe the loop properties we associate to each loop. We consider combinations of the following pragmas, based on Merlin’s optimizations, for the loop with iterator i: #pragma ACCEL parallel <factor=ufi>, #pragma ACCEL pipeline <II=IIi>, #pragma ACCEL tile and #pragma ACCEL cache <variable=a>.
We therefore associate to each loop i in the program a property vector that informs about the optimizations to be considered. We define $\vec{P}V_{i}$ as follows: $\vec{P}V_{i}:<ispipelined_{i},II_{i},uf_{i},tile_{i},TC^{min}_{i},TC^{max}_{i}>$ where we have: $ispipelined_{i}=1$ if the loop is pipelined, 0 otherwise; $II_{i}$ is the initiation interval, set to 1 by default; $uf_{i}$ is the parallelism/unroll factor, set to 1 by default (no #pragma ACCEL parallel pragma) and set to $TC^{max}_{i}$ if parallel is defined without a factor $uf_{i}$ specified. $tile_{i}$ is the TC of the innermost loop after strip mining. $TC^{min}_{i}$ is the minimal trip count of loop i, for any of its execution in the program. We also compute the maximal trip count over all executions. These values are computed using polyhedral analysis on the loops (Pouchet et al., 2013a). The pragma cache transfers above the loop i the data needed for the computation of this loop nest for the array a.
@@
Modeling Vitis optimizations  AMD/Xilinx Vitis will apply several optimizations automatically, such as auto-pipeline and auto-loop-flatten, some other optimizations when the user gives a compilation option such as tree reduction.
Only a loop with a constant $TC$, i.e., $TC_{max}=TC_{min}$ can be unrolled. The unroll pragma options allow to specify the unrolling factor, $uf_{i}$. When the factor is not specified, it implies that the factor is equal to the $TC$ of the loop. When a loop is pipelined, all innermost loops are automatically fully unrolled. Hence, we also propagate unrolling information, e.g., to mark a full loop nest for full unrolling if an outer loop is marked with #pragma ACCEL pipeline. The pipeline pragma options allow to specify the objective II the user wants to achieve. When II is not specified, it is automatically set to 1. In addition, Vitis will auto-pipeline with a target II of 1 the innermost loop which are not fully unrolled for each nested loop.
@@
Modeling Merlin optimizations  Finally, Merlin will also add automatic optimizations. It will explicitly strip-mine a loop when it is partially unrolled with the innermost loop having a $TC$ equal to the factor and the unrolled pragma applied to that loop. Similarly to Vitis, Merlin will auto-pipeline. Merlin also applies a program transformation for certain pragmas. When there are two perfectly nested and partially unrolled loops, Merlin swaps the two loops strip-mined (if legal), unrolled innermost and flatten and pipeline the two outermost loops. Further, Merlin will automatically transfer the data from off-chip to on-chip and cache on-chip with packing, with a maximum packing of 512 bits for our FPGA while computing if the footprint of the data fit on-chip by static analysis. The pragma tile allows Merlin to strip mine a loop and give the compiler the opportunity to transfer less/more data while respecting resource constraints. For our model we suppose an optimistic data transfer i.e., all memory transfers are done with a packing of 512 bits and each data are transferred once (perfect data reuse).
@@
4. Theoretical Latency and Resource Modeling
We now outline key elements of our analytical performance model, and the associated proofs this model computes a lower bound on latency under resource constraints. Full proofs cover several pages, and are available in the Appendix : we limit below to summarizing the hypothesis and overall proof approach.
@@ 4.1. Analytical Model Template
We take as input the summary AST of the program in constructor form, and the set of $\vec{P}V_{i}$ vectors for each loop, as defined in Sec. 3.1. Returning to our example in Sec. 3.1, this summary AST, capturing the loop structure and all statements, is $Loop_{i}(Loop_{j1}(S1),Loop_{j2}(S2,S3))$.
Let $I^{\vec{PV_{i}}}_{i}(X)$ be an operator that computes the latency of a region of a code represented by $X$, where $X$ is composed of loops and/or statements. This function is computed as follows: $I^{\vec{PV_{i}}}_{i}(X)=\lfloor II_{i}*(TC^{avg}_{i}/uf_{i}-pip_{i})\rfloor\odot X$ where $\odot$ is $+$ if the loop $i$ is pipelined ($pip_{i}=1$), $\times$ otherwise. Let $C_{i}^{\vec{PV_{i}}}(X_{1},...,X_{n})$ the operators which composed the region, defined as follows: $C_{i}^{\vec{PV_{i}}}(X_{1},...,X_{n})$ is $max_{i=1}^{N}(X_{i})$ if there is no dependency (WaR, RaW, WaW) between the statements of the regions and $\sum_{i=1}^{n}X_{i}$ between the statements which have at least one dependency. Finally, $SL^{\vec{PV_{i}}}_{i}(\vec{S_{k}})$ denotes a region of straight-line code (e.g., an inner-loop body). Intuitively, $SL$ will represent a lower bound on the latency of a code block such that, by the composition across all loops as per the template formula, the result remains a lower bound on the full program latency.
We build the analytical formula template by simple substitution of the operators in the summary AST by the following operators: (1) for every $Loop_{i}(\vec{X})$, replace it by $I^{\vec{P}V_{i}}_{i}(\vec{X})$ ; (2) for every list $\vec{I}$ containing the $I^{\vec{P}V_{i}}_{i}(\vec{X})$ of the step 1, replace it by $C^{\vec{PV_{i}}}_{i}(\vec{I})$; and (3) replace statement lists inside loop i $\vec{S}_{x}$ by $SL^{\vec{PV_{i}}}_{i}(\vec{S}_{x})$. That is for our example, we simply rewrite: $Loop_{i}(Loop_{j1}(S1),Loop_{j2}(S2,S3))$ into the following formula:
@@
4.2. Restrictions and Applicability
Our objective is to formulate a lower bound on the latency of a program, after HLS. We therefore include several restrictions:
  • The input program is a pure polyhedral program (Girbal et al., 2006), and its analysis (loop trip counts for every loop, all data dependencies (Feautrier, 1991)) is exact.
  • No HLS optimization shall change the number of operations in the computation: strength reduction, common sub-expression elimination, etc., shall either first be performed in the input program before analysis, or not be performed by the HLS toolchain.
  • The program does not contain ”useless” operations. A typical case is when a value is overwritten before being utilized, e.g., x = 42+51; x = 0;.
  • We only model DSP and BRAM resources for the considered kernel, ignoring all other resources.
  • We assume resources (DSP) sharing across different operations executing at the same cycle is not possible.
4.3. Latency Lower Bound
@@
Then, we can build a lower bound on the number of cycles a region $L$ may require to execute, under fixed resources, by simply taking the maximum between the weighted critical path and the work to execute normalized by the resources available.

### v2

@@ 2.4. Overview of NLP-DSE
NLP-DSE targets the (conservative) modeling of the performance and resources used by a design, such that arbitrary pragma configurations from Section 2.1 are applied on a regular, loop-based affine program. It deploys accurate static analysis to reason on the input program features, and a complex non-linear analytical performance model to drive the design space exploration. That is, NLP-DSE is a method for automatic pragma insertion that is specialized to affine programs. As demonstrated below, this specialization enables significantly better QoR and DSE time for affine programs than general-purpose DSE approaches such as AutoDSE.
@@
To make our approach feasible and maintain sufficient accuracy in analytical models, we focus on programs with static control flow that can be exactly captured using polyhedral structures (Feautrier, 1992). These affine, or polyhedral, programs can be analyzed to obtain the exact information about loop trip counts and dependencies, enabling more accurate performance predictors (Zuo et al., 2015b). The Affine MLIR dialect specializes in modeling such programs (Lattner et al., 2021; Ye et al., 2022).
Although NLP-DSE it can be used to compute pragmas without any DSE, the inherent limit of analytical models persist with NLP-DSE : as the implementation details of the back-end toolchain for HLS and synthesis may not be accurately captured by a model, DSE remains needed for best performance.
NLP-DSE enables the exploration of different parallelism and configuration spaces by constraining the level of parallelism, as detailed in Section 6. Our model is presented in Sec. 3, and we prove it is a performance lower bound in Sec. 4, an important feature to be able to prune designs during the search without the risk of losing performance. The associated NLP formulation is provided in Section 5.
The effectiveness of our framework is demonstrated in Table 3, comparing the performance and time-to-solution of NLP-DSE . We also display the result of the first synthesizable design produced, NLP-DSE -FS.
2mm Gemm Gramsch.
GF/s Time (mn) DSP (%) GF/s Time (mn) DSP (%) GF/s Time (mn) DSP (%)
Original 0.10 N/A 0 0.07 N/A 0 0.14 N/A 0
AutoDSE 0.41 1,870 14 68.91 1,345 10 0.95 819 1
NLP-DSE -FS 13.19 N/A 24 105.18 N/A 26 2.34 N/A 2
NLP-DSE 117.48 70 39 105.18 185 26 2.34 420 2
Imp. vs. AutoDSE 286x 26x 1.5x 7.2x 2.4x 1.9x
Table 3. Comparison of NLP-DSE , NLP-DSE -FS, which provides the result of the first synthesizable design, the source-to-source compiler Merlin without pragma insertion, and AutoDSE in terms of throughput (GF/s), DSE time (mn) and DSP utilization (%) for the kernels 2mm, Gemm, and Gramschmidt.
For Gemm and Gramsch the first design synthetizable has the best QoR of our DSE. For these two kernels, NLP-DSE implements better parallelism usage compared to AutoDSE. Specifically, our methodology successfully identified configurations with more balanced levels of parallelism. In contrast, AutoDSE failed to achieve the same level of parallelization within the fixed time for the DSE. On one hand, AutoDSE tends to explore first configurations with low levels of parallelism. On the other hand, it concurrently explores design spaces with excessively high levels of parallelism, leading to timeouts and unmet resource constraints. This discrepancy highlights the merits of seeding the DSE with configurations optimized for maximum parallelism, and systematically adjusts this level based on hardware directives and compiler expectations, as implemented in NLP-DSE .
For 2mm, the first design synthetizable allows us to have a better QoR vs. AutoDSE but our DSE shows its interest in finding a configuration 8.9 times faster than the first configuration found by the DSE. More detail can be found in Section 8.
It is noteworthy that our Design Space Exploration (DSE) approach, detailed in Section 6, deviates intentionally from AutoDSE. Unlike AutoDSE, which starts with a pragma-free configuration and gradually introduces pragmas, we begin with configurations characterized by the lowest theoretical latency, emphasizing high levels of parallelism. This deliberate departure from the conventional approach is further discussed in Section 6.
In Section 7, we present a comprehensive evaluation of our framework, demonstrating the improvements over AutoDSE that can be achieved by specializing to affine programs, in terms of design throughput and time-to-solutions across various benchmarks. The results indicate an average performance improvement of 5.69x and 17.24x in terms of DSE time and design throughput, respectively, with only a marginal decrease in throughput for 1 out of 47 benchmarks. Importantly, the time-to-solution of NLP-DSE it is consistently up to 30x faster than AutoDSE across all benchmarks.
@@ 3.1. Program Representation
We employ the code below as a running example with the pragma above the loops as AMD/Xilinx Merlin. For presentation simplicity, we assume each loop iterator in the program region has been renamed to a unique name, so that we can uniquely identify loops in the code by their iterator name.
@@ 5 S1(i,j1);
7 for(j2=lbj2(i);j2<=ubj2(i);j2++){
8 S2(i,j2);
9 S3(i,j2);
10 }
11}
The summary AST for this loop is simply built by creating one node per for loop and one per statement Sx, the body of a loop is made of loops and/or statements, and their nodes are children of said loop in the tree, listed in their syntactic order. For example, above, utilizing a constructor notation, it gives: $Loop_{i}(Loop_{j1}(S1),Loop_{j2}(S2,S3))$. Then, a simple rewrites of this tree using loop properties and composition operators will lead to the proposed analytical model, as outlined below. We first describe the loop properties we associate to each loop. We consider combinations of the following pragmas, based on Merlin’s optimizations, for the loop with iterator i:
  • #pragma ACCEL parallel <factor=ufi>
  • #pragma ACCEL pipeline <II=IIi>
  • #pragma ACCEL tile
  • #pragma ACCEL cache <variable=a>
We therefore associate to each loop i in the program a property vector that informs about the optimizations to be considered. We define $\vec{P}V_{i}$ as follows: $\vec{P}V_{i}:<ispipelined_{i},II_{i},uf_{i},tile_{i},TC^{min}_{i},TC^{max}_{i}>$ where we have: $ispipelined_{i}=1$ if the loop is pipelined, 0 otherwise; $II_{i}$ is the initiation interval, set to 1 by default; $uf_{i}$ is the parallelism/unroll factor, set to 1 by default (no #pragma ACCEL parallel pragma) and set to $TC^{max}_{i}$ if parallel is defined without a factor $uf_{i}$ specified. $tile_{i}$ is the TC of the innermost loop after strip mining. $TC^{min}_{i}$ is the minimal trip count of loop i, for any of its execution in the program. We also compute the maximal trip count over all executions. These values are computed using polyhedral analysis on the loops (Pouchet et al., 2013b). The pragma cache transfers above the loop i the data needed for the computation of this loop nest for the array a.
@@ 9 S3(i,j2);
Modeling Vitis optimizations
AMD/Xilinx Vitis will apply several optimizations automatically, such as auto-pipeline and auto-loop-flatten, some other optimizations when the user gives a compilation option such as tree reduction.
Only a loop with a constant $TC$, i.e., $TC_{max}=TC_{min}$ can be unrolled. The unroll pragma options allow to specify the unrolling factor, $uf_{i}$. When the factor is not specified, it implies that the factor is equal to the $TC$ of the loop. When a loop is pipelined, all innermost loops are automatically fully unrolled. Hence, we also propagate unrolling information, e.g., to mark a full loop nest for full unrolling if an outer loop is marked with #pragma ACCEL pipeline. The pipeline pragma options allow to specify the objective II the user wants to achieve. When II is not specified, it is automatically set to 1. In addition, Vitis will auto-pipeline with a target II of 1 the innermost loops which are not fully unrolled for each nested loop.
@@
Modeling Merlin optimizations
Finally, Merlin will also add automatic optimizations. It will explicitly strip-mine a loop when it is partially unrolled with the innermost loop having a $TC$ equal to the factor and the unrolled pragma applied to that loop. Similarly to Vitis, Merlin will auto-pipeline. Merlin also applies a program transformation for certain pragmas. When there are two perfectly nested and partially unrolled loops, Merlin swaps the two loops strip-mined (if legal), unrolled innermost and flatten and pipeline the two outermost loops. Further, Merlin will automatically transfer the data from off-chip to on-chip and cache on-chip with packing, with a maximum packing of 512 bits for our FPGA while computing if the footprint of the data fit on-chip by static analysis. The pragma tile allows Merlin to strip mine a loop and give the compiler the opportunity to transfer less/more data while respecting resource constraints. For our model we suppose an optimistic data transfer i.e., all memory transfers are done with a packing of 512 bits and each data are transferred once (perfect data reuse).
@@
4. Modeling of Latency and Resource Lower Bound
We present the foundational components of our analytical performance model and demonstrate how this model computes a lower bound on latency while adhering to resource constraints. The theorems and their accompanying proofs can be found in Appendix B.
@@ 4.1. Analytical Model Template
Consider the expression $I^{\vec{P}V_{i}}_{i}(X)$ for the loop $i$, where $X$ denotes a subpart of the program, and the operators are defined as follows:
$I^{\vec{PV}_{i}}_{i}(X)=\begin{cases}\left\lfloor II_{i}\cdot\left(\frac{TC^{avg}_{i}}{uf_{i}}-\text{ispip}_{i}\right)\right\rfloor+X&\text{if loop }i\text{ is pipelined}\\ &(\text{ispip}_{i}=1)\\ \\ \left\lfloor II_{i}\cdot\left(\frac{TC^{avg}_{i}}{uf_{i}}-\text{ispip}_{i}\right)\right\rfloor\cdot X&\text{otherwise }(\text{ispip}_{i}=0)\end{cases}$
For simplicity, we will represent the operation as $\odot$ to differentiate between the two different cases.
Let $C_{i}^{\vec{P}V_{i}}(X_{1},X_{2},\dots,X_{n})$ which compose the different sub part of a program under the loop $i$.
$C_{i}^{\vec{P}V_{i}}(X_{1},X_{2},\dots,X_{n})=\begin{cases}max(X_{1},X_{2},\dots,X_{n})&\text{ if }(X_{1},\dots,X_{n})\text{ do not have dependencies (WaR, RaW, WaW)}\\ \sum_{k=1}^{n}X_{k}&\text{otherwise}\end{cases}$
1S0: x1 += y;
2S1: x2 += x1 + z;
Listing 2: Example of code where two statements have dependencies
1S0: x1 += y;
2S1: x2 += z;
Listing 3: Example of code where two statements do not have dependency
In List 2, statements $S0$ and $S1$ exhibit dependencies, resulting in a program latency equal to the sum of their latencies, corresponding to $issequence=1$. Conversely, the statements in List 3 lack dependencies, allowing $S0$ and $S1$ to be executed concurrently. Therefore, the latency is determined by the maximum of $S0$ and $S1$ latencies, representing the case $inparallel=1$.
Finally, $SL^{\vec{P}V_{i}}_{i}(\vec{S}_{k})$ denotes a region of straight-line code (e.g., an inner-loop body). Intuitively, $SL$ will represent a lower bound on the latency of a code block such that by composition across all loops as per the template formula, the result remains a lower bound on the full program latency.
Given the summary AST of the program in constructor form, and the set of $\vec{P}V_{i}$ vectors for each loop, we build the analytical formula template as follows:
  1. (1)
    replace the $Loop_{i}(\vec{X})$ operator by $I^{\vec{P}V_{i}}_{i}(\vec{X})$ operator;
  2. (2)
    replace the $\vec{X}$ list by $C^{\vec{P}V_{i}}_{i}(\vec{X})$;
  3. (3)
    replace statement lists inside loop i $\vec{S}_{x}$ by $SL^{\vec{P}V_{i}}_{i}(\vec{S}_{x})$.
For our example, we can rewrite the expression:
$Loop_{i}(Loop_{j1}(S1),Loop_{j2}(S2,S3))$
as
@@
Upon substituting the definitions and assuming that the statements are independent, we obtain:
$\left\lfloor II_{i}\cdot\left(\frac{TC^{avg}_{i}}{uf_{i}}-\text{ispip}_{i}\right)\right\rfloor\odot\max\left(\left\lfloor II_{j1}\cdot\left(\frac{TC^{avg}_{j1}}{uf_{j1}}-\text{ispip}_{j1}\right)\right\rfloor\odot SL^{\vec{P}V_{j1}}_{j1}(S1),\left\lfloor II_{j2}\cdot\left(\frac{TC^{avg}_{j2}}{uf_{j2}}-\text{ispip}_{j2}\right)\right\rfloor\odot SL^{\vec{P}V_{j2}}_{j2}(S2,S3)\right)$
4.2. A Formal Model for Latency
Our objective is to formulate a lower bound on the latency of a program after HLS. We therefore have put several restrictions: we assume the input program is a polyhedral program, that is the control-flow is statically analyzable; all loops can be recognized and their trip count computed; and all array / memory accesses can be exactly modeled at compile-time. No conditional can occur in the program. While our approach may generalize beyond this class, we limit here to these strict assumptions.
To maintain a lower bound on latency by composition, we operate on a representation of (parts of the) program which is both schedule-independent and storage-independent: indeed, a lower bound on this representation is necessarily valid under any schedule and storage eventually implemented. We however require HLS to not change the count and type of operations. Furthermore, for lower bounding purposes, we assume unless stated otherwise $\forall i,~{}inparallel_{i}=1$. We will discuss in the next section a more realistic but compiler-dependent approach to set $inparallel_{i}$, based on dependence analysis.
We assume programs are made of affine loops, that are loops with statically computable control-flow, with loop bounds made only of intersection of affine expressions of surrounding loop iterators and program constants. We now assume loop bodies (i.e., statements surrounded by loops) have been translated to a list of statements, with at most a single operation (e.g., +,-,/,*) per statement. Operations are n-ary, that is they take $n\geq 0$ input scalar values as operand, and produce 0 or 1 output scalar value. A memory location can be loaded from (resp. stored to) an address stored in a scalar variable. This is often referred to as straight-line code . This normalization of the loop body facilitates the computation of live-in/live-out data for the code block, and the extraction of the computation graph. Note the region can be represented in Static Single Assignment form, to ensure different storage location for every assignment, facilitating the construction of the operation graph. In addition we require the input program to not contain useless operations which may be removed by the HLS toolchain e.g. by dead code elimination, as illustrated in Listing 4.
1int example(int y, int z){
2 int x;
3 S0: x = 12 + y; // dead-code elimination
4 S1: x = y + z;
5 return x;
6}
Listing 4: Example of Code Illustrating Dead Code Elimination: As Statement S1 writes to variable x, but the value of x assigned in Statement S0 is never used, the compiler will remove Statement S0.
The restriction can be summarized as:
  • The input program is a pure polyhedral program (Girbal et al., 2006), and its analysis (loop trip counts for every loop, all data dependences (Feautrier, 1992)) is exact.
  • No HLS optimization shall change the number of operations in the computation: strength reduction, common sub-expression elimination, etc. shall either first be performed in the input program before analysis, or not be performed by the HLS toolchain. The program also does not contain ”useless” operations that may be removed by the compiler.
  • We only model DSP and BRAM resources for the considered kernel, ignoring all other resources. We do not model LUT and FF resources, because from experience in the loop-based benchmarks we consider DSP and BRAM resources are the most constraining resources. Moreover, the estimation of LUTs and FFs is more tedious.
  • We assume resource (DSP) sharing across different operations executing at the same cycle is not possible.
@@ 4 S1: x = y + z;
We can compute the directed acyclic graph made of all statements (i.e., all n-ary operations), connecting all producer and consumer operations, to build the operation graph:
Definition 4.1 (Operation Graph).
Given a straight-line code region $R$ made of a list $L$ of statements $S\in L$, the operation graph OG is the directed graph <$<\{N,V_{I},root,V_{O}\},E>$ such that $\forall S_{i}\in L,~{}N_{S_{i}}\in N$; and for every operation with output $o$ and inputs $\vec{i}$ in $L$ $\forall S_{i},\forall i_{k}\in\vec{i}_{S_{i}},e_{i_{k},o}\in E$, $\forall S_{i}:(o_{S_{i}},{\vec{i}}_{S_{i}})\in L,S_{j}:(o_{S_{j}},{\vec{i}}_{S_{j}})\in L$ with $S_{i}\neq S_{j}$ then we have $E_{S_{i},S_{j}}\in E$ iff $o_{Si}\cap{\vec{i}}_{S_{j}}\neq\emptyset$. For every input (resp. output) in $S_{i}$ which is not matched with an output (resp. input) of another $S_{j}$ in $L$, create a node $V_{val}\in V_{I}$ (resp. $V_{O}$) for this input (resp. output) value. If $dim({\vec{i}}_{S_{i}})=0$ then an edge $e_{root,S_{i}}$ is added to $E$.
From this representation, we can easily define key properties to subsequently estimate the latency and area of this code region, such as its span, or critical path.
Definition 4.2 (Operation Graph critical path).
Given $OG^{L}$ an operation graph for region $L$. Its critical path $OG_{cp}$ is the longest of all the shortest paths between every pairs $(v_{i},v_{o})\in<\{V_{I},root\},V_{O}>$. Its length is noted #$\#OG_{cp}^{L}$.
On the graph on the Figure 1 the critical paths from $A[i]$ to $c,\forall i\in$ $[\![0,7]\!]$, have the same length.
4.2.1. Latency Lower Bound
We can build a simple a lower bound on the latency of an operation graph:
Theorem 4.3 (Lower bound on latency of an Operation Graph).
Given infinite resources, and assuming no operation nor memory movement can take less than one cycle to complete, the latency $LAT_{cp}^{L}\geq\#OG_{cp}^{L}$ is a lower bound on the minimal feasible latency to execute $L$.
We can then build a tighter lower bound on the number of cycles a region $L$ may take to execute, under fixed resources, by simply taking the maximum between the weighted span and the work to execute normalized by the resources available.